# When Every Token Counts: Optimal Segmentation for Low-Resource Language Models

### v2

@@ Abstract
Traditional greedy tokenization methods have been a critical step in Natural Language Processing (NLP), influencing how text is converted into tokens and directly impacting model performance. While subword tokenizers like Byte-Pair Encoding (BPE) are widely used, questions remain about their optimality across model scales and languages. In this work, we demonstrate through extensive experiments that an optimal BPE configuration significantly reduces token count compared to greedy segmentation, yielding improvements in token-saving percentages and performance benefits, particularly for smaller models. We evaluate tokenization performance across various intrinsic and extrinsic tasks, including generation and classification. Our findings suggest that compression-optimized tokenization strategies could provide substantial advantages for multilingual and low-resource (LR) language applications, highlighting a promising direction for further research and inclusive NLP.
@@
Bharath Raj S thanks: Equal Contribution, Garvit Suri , Vikrant Dewangan , Raghav Sonavane thanks: corresponding author, raghavsonavane@gmail.com Sprinklr AI
@@ 1 Introduction
Malaysian k ata c ak ları nd an from what they will say Future ‘acak’ (will) fragments into ‘c ak’, suffixes split into ‘ları’ (their) + ‘dan’ (from)
@@ 4.2 The Algorithm
Malaysian k ata c ak ları nd an kat acak ların dan 43 Preserves ‘acak’ (future) + ‘ların’ (their) + ‘dan’ (from) vs. ‘c ak ları nd’
@@ 5.2 Extrinsic Evaluation Tasks
Indonesian Emot Mei Silviana Saputri and Adriani (2018) Classification
@@ Appendix D Extrinsic Evaluation Tasks
  • Expanding to Indonesian, we employed two datasets from the indoNLU (Wilie et al., 2020; Mei Silviana Saputri and Adriani, 2018; Setya and Mahendra, 2018) collection : EmoT, which is an emotion classification dataset collected from Twitter consisting of tweets in Indonesian covering five emotion labels: anger, fear, happiness, love, and sadness; and WReTE, which is a textual entailment dataset constructed from Wikipedia revision history, containing pairs of sentences with binary semantic relations .

### v4

@@ Abstract
Traditional greedy tokenization methods have been a critical step in Natural Language Processing (NLP), influencing how text is converted into tokens and directly impacting model performance. While subword tokenizers like Byte-Pair Encoding (BPE) are widely used, questions remain about their optimality across model scales and languages. In this work, we demonstrate through extensive experiments that an optimal BPE configuration significantly reduces token count compared to greedy segmentation, yielding improvements in token-saving percentages and performance benefits, particularly for smaller models. We evaluate tokenization performance across various intrinsic and extrinsic tasks, including generation and classification. Our findings suggest that compression-optimized tokenization strategies could provide substantial advantages for multilingual and low-resource (LR) language applications, highlighting a promising direction for further research and inclusive NLP. The project page can be found here: https://vikr-182.github.io/loreslm/.
@@
Bharath Raj S , Garvit Suri , Vikrant Dewangan thanks: Equal Contribution, Raghav Sonavane thanks: corresponding author, raghavsonavane@gmail.com Sprinklr AI
@@ 4.2 The Algorithm
k ata c ak ları nd an kat acak ların dan 43 Preserves ‘acak’ (future) + ‘ların’ (their) + ‘dan’ (from) vs. ‘c ak ları nd’
@@ 5.2 Extrinsic Evaluation Tasks
Indonesian Emot Saputri et al. (2018) Classification
@@ Appendix D Extrinsic Evaluation Tasks
  • Expanding to Indonesian, we employed two datasets from the indoNLU (Wilie et al., 2020; Saputri et al., 2018; Setya and Mahendra, 2018) collection : EmoT, which is an emotion classification dataset collected from Twitter consisting of tweets in Indonesian covering five emotion labels: anger, fear, happiness, love, and sadness; and WReTE, which is a textual entailment dataset constructed from Wikipedia revision history, containing pairs of sentences with binary semantic relations .